OpenStudy (danjs):

The population of bacteria in a culture grows at a rate proportional to the number of bacteria present at time t [hrs.]. After 3 hours, 400 bacteria are present. After 10 hours, 2000 are present. What was the initial number of bacteria?

OpenStudy (danjs):

I'll give a hint - It is a linear first order ODE. \[\frac{ dP }{ dt } \alpha (P)\] so \[\frac{ dP }{ dt } = kP\]

OpenStudy (misty1212):

i will give you a bigger hint forget ode and fancy pants stuff after 7 hours it increases by a factor of \(2000\div 400=5\) use \[P=400\times 5^{\frac{t}{7}}\] you want to know what it was 3 hours earlier, plug in \(t=-3\)

OpenStudy (danjs):

Right, I am looking for the solution method to the general form of: \[\frac{ dy }{ dx } + p(x)*y = q(x)\] using an inegrating factor \[u(x) = e ^{\int\limits p(x)*dx}\]

ganeshie8 (ganeshie8):

\[\large 5^{t/7} = e^{(\ln 5/7)t}\]

ganeshie8 (ganeshie8):

why do you even need to solve a linear ODE ? your equation is separable right ?

ganeshie8 (ganeshie8):

\[\frac{ dP }{ dt } = kP\] \[\int\frac{ dP }{ P } = \int k~dt\]

ganeshie8 (ganeshie8):

solving gives you the same eqn used by misty

OpenStudy (misty1212):

why do you need to solve anything? you are given all the numbers you need in the question

ganeshie8 (ganeshie8):

@misty1212 the op is trying to derive that growth formula using calc it seems

OpenStudy (danjs):

Right, it is just a easy beginner problem as an intro to solving 1st order linear equations constant coefficients.

OpenStudy (danjs):

I just put it up for people to mess with.

OpenStudy (misty1212):

lol

ganeshie8 (ganeshie8):

its not a general linear ODE as you can see it is separable.. i think you need to add some gangsters that steal money from you every month or so to make it more general..

OpenStudy (danjs):

yes it is also separable i know..lol fine ill post one that is not separable, and needs to be solved with an integrating factor.

ganeshie8 (ganeshie8):

\[\frac{ dP }{ dt } = kP - 100t\] 100$ goes out eery time unit^ now it is no longer separable and you will have to use IF

ganeshie8 (ganeshie8):

*every

OpenStudy (danjs):

yeh, i was just making it easier. ill post a new one

ganeshie8 (ganeshie8):

I see.. y' = ky is linear but it is also separable and autonomous so we we solve it by separating variables adding some time factor on right hand side makes it impossible for separating variables

OpenStudy (mathmath333):

what is the answer

OpenStudy (mathmath333):

is it 252

OpenStudy (misty1212):

no 200

OpenStudy (danjs):

Just \[P(t) = P _{0}e ^{kt}\] Po is about 201 bacteria

OpenStudy (danjs):

\[k = \frac{ 1 }{ 7 }\ln(5)\]